# Solving Truly Massive Budgeted Monotonic POMDPs with Oracle-Guided Meta-Reinforcement Learning

### v1

@@ Abstract
Monotonic Partially Observable Markov Decision Processes (POMDPs), where the system state progressively decreases until a restorative action is performed, can be used to model sequential repair problems effectively. This paper considers the problem of solving budget-constrained multi-component monotonic POMDPs, where a finite budget limits the maximal number of restorative actions. For a large number of components, solving such a POMDP using current methods is computationally intractable due to the exponential growth in the state space with an increasing number of components. To address this challenge, we propose a two-step approach. Since the individual components of a budget-constrained multi-component monotonic POMDP are only connected via the shared budget, we first approximate the optimal budget allocation among these components using an approximation of each component POMDP’s optimal value function which is obtained through a random forest model. Subsequently, we introduce an oracle-guided meta-trained Proximal Policy Optimization (PPO) algorithm to solve each of the independent budget-constrained single-component monotonic POMDPs. The oracle policy is obtained by performing value iteration on the corresponding monotonic Markov Decision Process (MDP). This two-step method provides scalability in solving truly massive multi-component monotonic POMDPs. To demonstrate the efficacy of our approach, we consider a real-world maintenance scenario that involves inspection and repair of an administrative building by a team of agents within a maintenance budget. Our results show that the proposed method significantly improves average component survival times compared to baseline policies, thereby highlighting its potential for practical applications in large-scale maintenance problems. Finally, we perform a computational complexity analysis for a varying number of components to show the scalability of the proposed approach.
@@ 1 Introduction
Partially Observable Markov Decision Processes (POMDPs) provide an efficient framework for modeling a wide range of real-world sequential decision-making problems (Cassandra 1998; Bravo, Leiras, and Cyrino Oliveira 2019). Numerous methods have been developed to compute optimal policies for POMDPs, including Monte-Carlo
@@
methods (Katt, Oliehoek, and Amato 2017), reinforcement learning approaches (Singh et al. 2021), and various approximation techniques (Kearns, Mansour, and Ng 1999). One notable application of POMDPs is in infrastructure maintenance or sequential repair problems, where the system state represents the condition of infrastructure components, which is only partially observable (Bhattacharya et al. 2021). Over time, the state of these components stochastically deteriorates unless a restorative action, such as a repair, is performed. Figure 1 shows this stochastic decrease and also shows the probability distribution of the condition of a sample infrastructure component for different time steps. The state evolution of the POMDP, thus, exhibits monotonic properties (Miehling and Teneketzis 2020) and we define such POMDPs as monotonic POMDPs. Previously, Bhattacharya et al. (2020) proposed an optimal inspection and repair scheduling policy for a pipeline, which is a single-component POMDP. However, infrastructure systems naturally have multiple components (Daulat et al. 2024).
@@
In this paper, we address the challenge of computing approximately optimal policies for budget-constrained multi-component monotonic POMDPs. Each component POMDP operates independently in terms of transition probabilities, but they are collectively constrained by the shared budget. There has been substantial work on solving budget constrained POMDPs (Lee et al. 2018; Undurti and How 2010; Khonji, Jasour, and Williams 2019). However, the complexity of these algorithms is often exponential in the number of states of the POMDP, and consequently would be exponential in the number of components for a multi-component POMDP. Thus, they become computationally intractable for multi-component POMDPs with a large number of components. In Vora et al. (2023), the authors propose a welfare-maximization method for solving budget-constrained multi-component POMDPs. However, the method in that paper requires generating optimal policies for multiple budget values for every component POMDP to get the optimal budget allocation, i.e., optimal distribution of the shared budget among the components POMDPs. Hence, it cannot be scaled to a large number of components.
To address this scalability issue, we propose a computationally efficient meta-reinforcement learning approach for solving budget-constrained multi-component monotonic POMDPs. First, we approximate the optimal distribution of budget among the individual component POMDPs using a random forest budget allocation strategy, thereby decoupling the problem into independent components. The random forest model is trained to approximate the value function of a component POMDP as a function of the budget, leveraging the concavity of the value function to frame the budget allocation as a concave maximization problem, which guarantees a computationally feasible solution. Subsequently, we apply the oracle policy-guided meta-trained Proximal Policy Optimization (PPO) algorithm, trained across various components and budget values, to compute the approximately optimal policy for each component-budget pair. The oracle policy for the meta-PPO is derived from value iteration applied to the corresponding component MDP. To validate the efficacy of our proposed approach, we examine a maintenance scenario involving an administrative building with 1000 infrastructure components and a fixed maintenance budget. We compare the performance of our approach, in terms of the survival time of the components, against the oracle policy, a currently-used baseline policy, and a vanilla meta-PPO algorithm. We also perform a computational complexity analysis of our algorithm for varying number of components and show that the proposed approach is linear in the number of components.
In summary, the key contributions of this paper are:
@@
    We develop an oracle policy-guided meta-PPO algorithm for solving budget-constrained single-component monotonic POMDPs.
@@
    We introduce a random forest budget allocation strategy for optimally distributing the budget among the component POMDPs of a budget-constrained multi-component POMDP.
@@
    We demonstrate the efficacy of our approach through a large-scale real-world infrastructure maintenance scenario with 1000 infrastructure components.
@@
    We show the scalability of our algorithm through a computational complexity analysis for varying number of components.
@@ 2.1 Partially Observable Markov Decision Processes
A discrete-time finite-horizon Partially Observable Markov Decision Process (POMDP) (Cassandra, Kaelbling, and Littman 1994) $M$ is defined by the 7-tuple $(\mathcal{S},A,T,\Omega,O,R,H)$, which denotes the state space, action space, state transition function, observation space, observation function, reward function and planning horizon, respectively. In a POMDP, the agent does not have direct access to the true state of the environment. Instead, the agent may maintain a belief state, representing a probability distribution over $\mathcal{S}$. This belief is updated based on the received observation using Bayes’ rule (Araya et al. 2010).
@@ 2.2 POMDP Solution Methods
Computing optimal policies for a POMDP is generally PSPACE-complete (Mundhenk et al. 2000; Vlassis, Littman, and Barber 2012). Thus, to address the computational intractability of solving POMDPs, various approximation methods have been widely used (Poupart and Boutilier 2002; Pineau et al. 2003; Roy, Gordon, and Thrun 2005). Several reinforcement learning approaches have also been developed for computing approximate POMDP solutions (Azizzadenesheli, Lazaric, and Anandkumar 2016; Igl et al. 2018). However these methods become computationally intractable when faced with the high dimensionality and shared resource constraints of budget-constrained multi-component monotonic POMDPs such as those considered in this paper.
@@ 2.3 Consumption MDPs and Budgeted POMDPs
The integration of budget or resource constraints into Markov Decision Processes (MDPs) has been previously studied under the frameworks of Consumption MDPs (Blahoudek et al. 2020) and Budgeted POMDPs (Vora et al. 2023). However, the algorithm proposed in Blahoudek et al. (2020) assumes full observability of the state and hence cannot be applied to budget-constrained POMDPs. A solution for budget-constrained multi-component POMDPs is presented in Vora et al. (2023). However, the method in this paper requires repeated computations of optimal policies for different budget values for all component POMDPs and hence is not scalable to a budget-constrained multi-component POMDP with a large number of components.
@@ 3 Problem Formulation
$T^{i}(s^{i}_{k},a^{i}_{k},s^{i}_{k+1})=\begin{cases}p_{1}^{i}(s^{i}_{k},a^{i}_{k},s^{i}_{k+1}),&\text{if }a^{i}_{k}=m^{i}\text{ and }\\ &s^{i}_{k}\leq s^{i}_{k+1}\leq\bar{s},\\ p_{2}^{i}(s^{i}_{k},a^{i}_{k},s^{i}_{k+1}),&\text{if }a^{i}_{k}\in\{d^{i},q^{i}\}\\ &\text{and }s^{i}_{k+1}\leq s^{i}_{k},\\ 1,&\text{if }a^{i}_{k}\in A^{i}\text{ and }\\ &s^{i}_{k+1}=0=s^{i}_{k},\\ 0,&\text{otherwise}.\end{cases}$
@@ 3.1 Problem Statement
The primary objective of this paper is to determine a policy $\pi^{*}$ for this multi-component monotonic POMDP over a horizon $H$, that maximizes the sum of expectations of the individual times before reaching the absorbing state for each component, while adhering to the total budget $B$. We denote this maximal time $k$ by $T_{max}=\sum_{i=1}^{n}T^{i}_{max}$, where $T^{i}_{max}$ denotes the corresponding maximal time for component $i$. Mathematically, the problem can be formulated as:
@@ 4 Solution Approach
In this section, we present our methodology for solving a budget-constrained multi-component monotonic POMDP. Figure 2 presents an architectural overview of our proposed approach. We first input information about all $n$ component POMDPs into a pre-trained random forest regressor to get the $T^{i}_{max}$ for each component $i$ as a function of the budget allocated to the component. Next, we propose an appropriate allocation of the shared budget $B$ among the component POMDPs by solving a constrained maximization problem. We then compute the oracle policy for each component POMDP and budget pair through value iteration applied to the corresponding component MDP. Finally, using these oracle policies and a meta-trained reinforcement learning (RL) agent, we solve each component POMDP with respect to the allocated budget and consequently propose a policy for the multi-component POMDP. Note that an alternate allocation strategy could involve redistributing the budget at every time step during planning. However, such a method would be computationally more expensive, due to the repeated computation of the allocation, as compared to our proposed a priori budget distribution approach.
4.1 Oracle-Guided RL for Budget-Constrained Monotonic POMDPs
To derive the optimal policy for a budget-constrained single-component monotonic POMDP, we propose an oracle policy-guided reinforcement learning algorithm. Adherence to the budget constraints is achieved using the budget-constrained POMDP (bPOMDP) framework proposed in Vora et al. (2023). In a bPOMDP, the state at each time step includes an additional fully observable component representing the cost incurred up to that point.

### v2

@@ Abstract
Many real-world decision problems, ranging from asset-maintenance scheduling to portfolio rebalancing, can be naturally modelled as budget-constrained multi-component monotonic Partially Observable Markov Decision Processes (POMDPs): each component’s latent state degrades stochastically until an expensive restorative action is taken, while all assets share a fixed intervention budget. For a large numbers of assets, deriving an optimal policy for this joint POMDP is computationally intractable. To tackle this challenge, we prove that the value function of the associated belief-MDP is budget-concave, which allows an efficient two-step approach to finding a near-optimal policy. First, we approximate the optimal cross-component budget split via a random-forest surrogate of each single-component value function. Second, we solve each resulting budget-constrained single-component POMDP with an oracle-guided meta-trained Proximal Policy Optimization (PPO) policy: value-iteration on the fully observable counterpart yields an oracle that shapes the PPO update and greatly accelerates learning. We validate our method through experiments in two disparate domains: (i) preventive maintenance for a large-scale building infrastructure containing 1,000 components, and (ii) portfolio risk management under debit-only loss-budget constraints, where each asset’s latent budget depletes with market losses and can only be replenished through costly recapitalization. Results show that our method consistently achieves longer component survival times and enhanced portfolio viability than both baseline heuristics and vanilla PPO. Furthermore, our approach maintains linear scalability in solution time with respect to the number of components.
@@ 1 Introduction
Partially Observable Markov Decision Processes (POMDPs) offer a principled framework for sequential decision making under uncertainty regarding the true state of the system  Cassandra (1998); Bravo et al. (2019). Solving POMDPs is computationally challenging, leading to the development of various solvers, including Monte-Carlo tree search Katt et al. (2017), reinforcement-learning variants Singh et al. (2021), and diverse approximation schemes Kearns et al. (1999).
Many application domains share a monotonic structure, where the latent state of individual components degrades stochastically over time unless a costly restorative action is taken. Canonical examples include online advertising  Boutilier & Lu (2016), inventory replenishment Shin & Lee (2015), and sequential repair or maintenance scheduling for physical assets Miehling & Teneketzis (2020); Bhattacharya et al. (2021).
@@
Figure 1 shows this stochastic decline and the probability distribution of a sample component’s condition at successive time steps. While prior work, such as Bhattacharya et al. (2020), has addressed optimal policies for single-component systems, real-world systems—from building portfolios to exchange-traded-fund (ETF) baskets—naturally involve many such components Daulat et al. (2024).
@@
In this paper, we address the challenge of computing approximately optimal policies for budget-constrained multi-component monotonic POMDPs. We assume that each component POMDP operates independently in terms of transition probabilities, but they are collectively constrained by the shared budget. Substantial work has been done to solve budget-constrained POMDPs Lee et al. (2018); Undurti & How (2010); Khonji et al. (2019). However, the complexity of these algorithms is often exponential in the number of states of a single POMDP. For a multi-component POMDP, where the overall state space is the Cartesian product of individual component state spaces, this complexity consequently becomes exponential in the number of components. Thus, these methods become computationally intractable for multi-component POMDPs with a large number of components. A key challenge in solving budget-constrained multi-component POMDPs is how to optimally allocate the shared budget across the multiple components. In Vora et al. (2023), the authors propose a welfare-maximization method for solving budget-constrained multi-component POMDPs. However, the method in that paper requires generating optimal policies for multiple budget values for every component POMDP to get the optimal budget allocation. Hence, it cannot be scaled to a large number of components.
Our insight.
The primary computational bottleneck in solving budget-constrained multi-component POMDPs is the coupling induced by the shared budget. If that budget could be split a-priori, the joint POMDP would factor into $n$ independent, single-component problems solvable in parallel. To enable this decomposition, we prove that the optimal value function of a single monotonic POMDP is concave in its allocated budget. This budget-concavity lets us decouple first, optimise second:
  1. (1)
    Budget allocation. We maximize a concave surrogate of the value function, estimated with a random-forest regressor, to distribute the global budget across components; and
  2. (2)
    Component policies. With budgets fixed, we learn a near-optimal policy for each component–budget pair using an oracle-guided, meta-trained Proximal Policy Optimization (PPO) agent, where the oracle is obtained by value iteration on the fully observable counterpart.
The result is a scalable solution whose runtime grows linearly with the number of components while retaining strong performance guarantees.
Contributions.
@@
    Theory. We prove budget-concavity of the optimal value function for monotonic POMDPs. While prior works implicitly assume and use this budget-concavity, our work provides the first general structural guarantee that formally links budget availability to expected return.
@@
    Algorithms. We introduce (i) a random-forest budget-allocation module that exploits concavity for fast global optimization, and (ii) an oracle-guided meta-PPO solver for each single-component POMDP.
@@
    Empirical evidence. On two domains—preventive maintenance of a 1000-component building and portfolio loss-budget management with recapitalization—we outperform baseline heuristics and vanilla PPO, whilst solution time of the proposed approach scales linearly in the number of components.
@@
    Complexity analysis. We provide a detailed runtime study confirming linear growth in wall-clock time as components increase from $n=10$ to $n=1000$.
The remainder of the paper is organized as follows. Section 2 surveys related work on budget-constrained POMDPs and large-scale maintenance or portfolio problems. Section 3 formalises the budget-constrained multi-component monotonic POMDP. Section 4 presents our solution pipeline: (i) Subsection 4.1 proves budget–concavity of the single-component value function; (ii) Subsection 4.2 exploits this structure to allocate the global budget via a random-forest surrogate; and (iii) Subsection 4.3 derives an oracle-guided meta-PPO policy for each component and composes them into the overall controller. Section 5 reports empirical results on infrastructure maintenance and financial loss-budget management, and Section 6 concludes with key findings and future directions.
@@ 2.1 Partially Observable Markov Decision Processes
A discrete-time finite-horizon Partially Observable Markov Decision Process (POMDP) Cassandra et al. (1994) $M$ is defined by the 7-tuple $(\mathcal{S},A,T,\Omega,O,R,H)$, which denotes the state space, action space, state transition function, observation space, observation function, reward function and planning horizon, respectively. In a POMDP, the agent does not have direct access to the true state of the environment. Instead, the agent may maintain a belief state, representing a probability distribution over $\mathcal{S}$. This belief is updated based on the received observation using Bayes’ rule Araya et al. (2010).
@@ 2.2 POMDP Solution Methods
Computing optimal policies for a POMDP is generally PSPACE-complete Mundhenk et al. (2000); Vlassis et al. (2012). Thus, to address the computational intractability of solving POMDPs, various approximation methods have been widely used Poupart & Boutilier (2002); Pineau et al. (2003); Roy et al. (2005). Several reinforcement learning approaches have also been developed for computing approximate POMDP solutions Azizzadenesheli et al. (2016); Igl et al. (2018). However, these methods become computationally intractable when faced with the high dimensionality and shared resource constraints of budget-constrained multi-component monotonic POMDPs such as those considered in this paper.
@@ 2.3 Consumption MDPs and Budgeted POMDPs
The integration of budget or resource constraints into Markov Decision Processes (MDPs) has been previously studied under the frameworks of Consumption MDPs Blahoudek et al. (2020) and Budgeted POMDPs Vora et al. (2023). However, the algorithm proposed in Blahoudek et al. (2020) assumes full observability of the state and hence cannot be applied to budget-constrained POMDPs. A solution for budget-constrained multi-component POMDPs is presented in Vora et al. (2023). However, the method in this paper requires repeated computations of optimal policies for different budget values for all component POMDPs and hence is not scalable to a budget-constrained multi-component POMDP with a large number of components.
@@ 3 Problem Formulation
[…]
@@ 3.1 Problem Statement
The primary objective of this paper is to determine a policy $\pi^{*}$ for this multi-component monotonic POMDP over a horizon $H$, that maximizes the sum of expectations of the individual times before reaching the absorbing state for each component, while adhering to the total budget $B$. We denote this maximal time $k$ by $T_{max}=\sum_{i=1}^{n}T^{i}_{max}$, where $T^{i}_{max}$ denotes the corresponding maximal time for component ii. Mathematically, the problem can be formulated as:
@@ 4 Solution Approach
In this section, we present our methodology for solving a budget-constrained multi-component monotonic POMDP. Figure 2 presents an architectural overview of our proposed approach. The key idea is to decouple first, optimize second. Allocating the shared budget as a first step of planning shrinks the original large joint POMDP into $n$ independent single-component POMDPs. Each of these single-component POMDPs now operates with its own fixed budget cap, which is determined by the initial allocation. This transformation converts a problem that is intractable for $n\!\gg\!1$ into $n$ modest ones that can be solved in parallel. We organize the section accordingly:
  • Structural Result: Budget Concavity (Section 4.1): We prove that each component’s value function is concave in its budget.
  • Stage 1: Budget Allocation (Section 4.2): Leveraging concavity, we fit a random-forest surrogate of the value function and solve a tractable concave maximization problem to distribute the shared global budget across components.
  • Stage 2: Oracle-Guided Meta-PPO (Section 4.3): With budgets fixed, we learn near-optimal policies for each component–budget pair (with respect to that component’s allocated budget and local POMDP) using an oracle-guided, meta-trained PPO agent, then compose these into the overall multi-component policy.
Note that an alternate allocation strategy could involve redistributing the budget at every time step during planning. However, such a method would be computationally more expensive than our proposed approach due to the repeated computation of the allocation.
4.1 Budget–Concavity of the Value Function
We first show that the optimal value function of a single-component belief-MDP—derived from a monotonic POMDP with random, nonnegative action costs—is concave in the budget variable $B$ for any finite planning horizon $H\geq 0$.
Setting
We consider a POMDP defined by the tuple $\langle S,A,T,R,\Omega,O,\gamma,\mathcal{C}\rangle$, where:
  • $S$ is a finite state space and AA is a finite action space, as defined in Section 3.
  • $T(s^{\prime}|s,a)$ is the transition kernel (1); $R(s,a)$ is the reward function.
  • $\Omega$ is the observation space; $O(o|s^{\prime},a)$ is the observation model as defined in Section 3.
  • $\gamma\in[0,1)$ is the discount factor.
  • $\mathcal{C}(s,a)$ is the distribution of a random, nonnegative cost incurred by taking action $a$ in state $s$. The specific cost, $c$, is drawn from this distribution.
Following standard practice in POMDP literature Cassandra (1998), we reformulate this POMDP as a belief-MDP with state $(b,B)$, where $b\in\Delta(S)$ is the belief (posterior distribution over hidden states), and $B\geq 0$ is a remaining budget. The reward at belief $b$ under action $a$ is:
$\rho(b,a):=\sum_{s\in S}b(s)R(s,a),$
and the budget evolves as $B\mapsto B-c$, where $c$ is a realization from the random cost $C(s,a)$ under the current belief. To prove the budget-concavity of the value function for this belief-MDP, we first establish two foundational properties concerning concavity. These lemmas demonstrate how concavity is preserved under common mathematical operations relevant to dynamic programming.
Lemma 1 (Concavity under Affine Shift).
If $f(x)$ is concave on an interval II, then f(xl)f(x-l) is concave on the interval $\{y\mid y=x+l,x\in I\}$ for any constant $l$.
Proof.
This lemma is a standard result in convex analysis Boyd & Vandenberghe (2004). ∎
Lemma 2 (Expectation Preserves Concavity).
Let $f(B,\xi)$ be concave in $B$ for every realization $\xi$. If $\xi$ is a random variable following an arbitrary probability distribution, then $\mathbb{E}_{\xi}[f(B,\xi)]$ is also concave in $B$.
Proof.
Fix any $B_{1},B_{2}\in\mathbb{R}$ and any $\lambda\in[0,1]$. Let $g(B)=\mathbb{E}_{\xi}[f(B,\xi)]$. Then
$g(\lambda B_{1}+(1-\lambda)B_{2})=\mathbb{E}_{\xi}\left[f\big{(}\lambda B_{1}+(1-\lambda)B_{2},\xi\big{)}\right].$
Since $f(B,\xi)$ is concave in $B$ for each $\xi$, we have
$f\big{(}\lambda B_{1}+(1-\lambda)B_{2},\xi\big{)}\geq\lambda f(B_{1},\xi)+(1-\lambda)f(B_{2},\xi)$
for all $\xi$. Taking expectations on both sides yields
$\mathbb{E}_{\xi}\left[f(\lambda B_{1}+(1-\lambda)B_{2},\xi)\right]\geq\lambda\mathbb{E}_{\xi}\left[f(B_{1},\xi)\right]+(1-\lambda)\mathbb{E}_{\xi}\left[f(B_{2},\xi)\right],$
that is,
$g(\lambda B_{1}+(1-\lambda)B_{2})\geq\lambda g(B_{1})+(1-\lambda)g(B_{2}).$
Thus, $g$ is concave in $B$. ∎
Having established these fundamental properties regarding the preservation of concavity under affine shifts and expectations, we will now prove that the optimal value function of a monotonic POMDP is concave with respect to the available budget.
Theorem 3 (Budget Concavity).
For any fixed belief $b\in\Delta(S)$ and horizon $H\geq 0$, the value function $V_{H}(b,B)$ is concave in $B$ on $0,\infty)$.
Proof.
We proceed by mathematical induction on $H$.
Base Case ($H=0$). At horizon zero, there are no rewards:
$V_{0}(b,B)=0\quad\text{for all }b\in\Delta(S),\ B\geq 0.$
Function $V_{0}$ is thus trivially concave.
Inductive Hypothesis. Suppose that for some $H\geq 0$, the function $V_{H}(b,B)$ is concave in $B$ for every belief $b$.
Inductive Step. We aim to prove that $V_{H+1}(b,B)$ is concave in $B$ for all $b$. The Bellman equation in the belief-MDP is:
$V_{H+1}(b,B)=\max_{a\in A}\left\{\rho(b,a)+\gamma\,\mathbb{E}_{o,c\mid b,a}\left[V_{H}(b^{\prime},B-c)\right]\right\},$
where $b^{\prime}$ is the updated belief after taking action $a$ and observing $o$. The expectation $\mathbb{E}_{o,c\mid b,a}$ is taken over the random observation $o$ and cost $c$ given the current belief $b$ and chosen action $a$.
Define the inner expectation as:
$g(a,b,B):=\mathbb{E}_{o,c\mid b,a}\left[V_{H}(b^{\prime},B-c)\right].$
Apply Lemma 1 and inductive hypothesis to assert that $B\mapsto V_{H}(b^{\prime},B-c)$ is concave for each $(o,c)$. Then Lemma 2 implies that $g(B)$, being the expectation over such functions, is also concave.
Therefore, the $Q$-value
$Q_{H+1}(b,B,a):=\rho(b,a)+\gamma g(B)$
is concave in $B$ for each $a$.
Finally, the value function is
$V_{H+1}(b,B)=\max_{a\in A}Q_{H+1}(b,B,a),$
which is the pointwise maximum of finitely many concave functions, and hence concave itself. ∎
4.1.1 Relating $\mathbb{E}[T_{\max}]$ to the Value Function
We proved the budget-concavity of the value function $V_{H}(s,B)$ in Theorem 3. In our problem setting (2), however, we aim to maximize the expected time to failure $E[T_{\max}]$. Specifically, in many practical applications such as preventive maintenance or portfolio management, the objective can be naturally framed as maximizing the expected time until a critical failure occurs or a budget is exhausted. We now show how the concavity property extends to $E[T_{\max}]$, which serves as the objective function for our initial budget allocation stage.
Lemma 4 (Expected-time equivalence).
Consider the reward function
$R(s,a)=\begin{cases}1,&s\neq 0,\\ 0,&s=0,\end{cases}$ (4)
. Let $V(s,B)$ be the corresponding optimal value function. Denote by $\mathbb{E}[T_{\max}(B)]$ the expected time to reach the absorbing state $s=0$ under the optimal budget-feasible policy. Then
$V(s,B)=\mathbb{E}[T_{\max}(B)].$
Proof.
Under reward scheme (4) each non-absorbing step contributes exactly 11 to the return; steps in state 0 contribute 0. Hence, for any budget-feasible policy $\pi$,
$\text{Total reward}=\mathbb{E}\Bigl{[}\sum_{t=0}^{H}\mathbf{1}\{s_{t}\neq 0\}\Bigr{]}=\mathbb{E}[T_{\max}(\pi)].$
Taking the maximum over all budget-feasible policies yields $V(s,B)=\mathbb{E}[T_{\max}(B)]$. ∎
Corollary 5.
Because $B\mapsto V(s,B)$ is concave by Theorem 3, the expected absorption time $\mathbb{E}[T_{\max}(B)]$ is also concave in the allocated budget $B$.
4.2 Random Forest Approach for Optimal Budget Allocation
By Corollary 5, the expected maximal survival time $\mathbb{E}[T_{\max}(B)]$ is a concave function of the budget allocated to a single component. This structural property lets us treat budget splitting across $n$ components as a concave maximization problem—one that is both tractable and amenable to surrogate modeling. Each component evolves independently but competes for the shared budget, rendering the components weakly coupled. While reinforcement learning algorithms have made significant advances, they often face challenges when scaling to the extremely large state and action spaces characteristic of multi-component systems Sutton & Barto (2018). To address this scalability issue, our remedy is an a-priori budget distribution that decouples the system. Concretely, for component $i$ we approximate the concave map $B\mapsto\mathbb{E}[T_{\max}^{i}(B)]$ by the exponential surrogate
$\widetilde{T}^{i}_{\max}(B)\;=\;\alpha^{i}\,e^{\beta^{i}B}+\gamma^{i},$ (5)
where $(\alpha^{i},\beta^{i},\gamma^{i})$ are constants. While many other concave functions could be used to model $\tilde{T}^{i}_{\max}$, we empirically observe that the exponential function provides a good fit for the data (see Appendix A). We use a random forest regressor Breiman (2001) to estimate the parameters of this exponential function. The training dataset for this model is obtained via non-linear least squares regression on multiple $(\mathbb{E}[T_{\max}],b)$ pairs for various budget-constrained single-component monotonic POMDPs. The input to this model includes specific statistics related to the POMDP’s transition function, which are the expected time to reach state 0 without repairs, $\mathbb{E}[T]$, and the variance of this expected time, $\sigma^{2}_{\mathbb{E}[T]}$, as well as the various actions costs.
Let $b^{i}$ denote the budget assigned to component $i$ and $\widetilde{T}^{i}_{\max}$ its surrogate survival time. The allocation problem becomes
$\max_{b^{1:n}}$ $\sum_{i=1}^{n}\widetilde{T}^{i}_{\max}(b^{i})$ (6)
s.t. $\sum_{i=1}^{n}b^{i}\;\leq\;B,\qquad b^{i}\;\geq 0\;\;\forall i,$
a concave maximization with linear constraints. Because each surrogate in (5) is concave, the problem is globally tractable and we solve it with off-the-shelf convex optimizers. Solving (6) yields the approximately optimal budget allocation among the individual components. The next subsection shows how an oracle-guided meta-PPO agent learns the individual component policies given this budget allocation.
4.3 Oracle-Guided RL for a Budget-Constrained Single Component
Given the per-component budgets $b^{i}$ obtained in Section 4.2, we now derive a near-optimal control policy for each single-component budget-constrained monotonic POMDP. We adopt the budget-augmented POMDP (bPOMDP) formalism of Vora et al. (2023), in which the state includes an additional, fully-observable coordinate that tracks cumulative cost.